OpenStudy (anonymous):

solve on the interval (0, 2pi): 3 csc x - 1 = 2/3

OpenStudy (driftracer305):

k so try isolating csc x

OpenStudy (anonymous):

csc x = 2/3 + 1?

OpenStudy (driftracer305):

lol........ where did the 3 from csc side go?

OpenStudy (anonymous):

oh crap! the initial equation is "3 csc x - 1 = 2"! :/ i had already made an attempt to isolate the beginning and confused myself

OpenStudy (driftracer305):

k i will show u....... try following........ if u have question.... pls interrupt me in the middle ok?

OpenStudy (anonymous):

right

OpenStudy (driftracer305):

k so first.......... csc x - 1 = (2/3)/3 ............ the 3 comes here to divide 2/3............. so it will be csc x -1 = 2/9.............. all good?

OpenStudy (anonymous):

no wait we don't divide (2/3) by three because the initial answer was 2 and the question i wrote wasn't supposed to be (2/3), just 3

OpenStudy (anonymous):

DANG IT I MEANT JUST 2

OpenStudy (driftracer305):

so the question was......... 3 csc x -1 = 2.....?

OpenStudy (anonymous):

yes

OpenStudy (driftracer305):

oh wow........... ok neva mind......... k so this time same..... divide..... so we get csc x - 1 = 2/3........... agree?

OpenStudy (anonymous):

agree

OpenStudy (driftracer305):

k now.......... bring the 1 to the other side......... so we get csc x = 2/3 + 1.... which equals csc x = 5/3........ agree?

OpenStudy (anonymous):

because 1 is a whole number so we make it 3/3 to add.... gotcha so far.

OpenStudy (driftracer305):

yup.......... so in anyway... the addition is 5/3........ all good right?

OpenStudy (anonymous):

yes

OpenStudy (driftracer305):

k now all u do is find x....... so now here to find x we take inverse of csc....... this makes the equation look like csc^-1 (5/3)........... and after solving this u get x................. all done......!!

OpenStudy (driftracer305):

use calculator and tell me what u get......

OpenStudy (driftracer305):

@giniasal ....... u there?

OpenStudy (anonymous):

yah i just had to grab my calculator

OpenStudy (driftracer305):

k

OpenStudy (anonymous):

ehm... i think i did this wrong... to get csc on a calculator is 1/sin so..if i'm doing inverse i'd put 1/sin^-1 (5/3)?

OpenStudy (driftracer305):

no actually csc^-1 is actually 1/csc.......... so u wuld have 1/(1/sin)...... so sin(5/3)

OpenStudy (driftracer305):

but @giniasal ........ to be on the safe side....... just use csc^-1

OpenStudy (driftracer305):

k do u have choices?

OpenStudy (driftracer305):

answer choices?

OpenStudy (anonymous):

yah man, there's no csc option on my calculator (TI-83)

OpenStudy (driftracer305):

ohh......... ok no prb........ use this..... http://www.wolframalpha.com/input/?i=csc%5E-1%285%2F3%29

OpenStudy (driftracer305):

ur answer is the degree part..... which is 36.81....

OpenStudy (anonymous):

wow thanks so much!

OpenStudy (driftracer305):

sure....... anytime... glad to help...... :D :D

OpenStudy (anonymous):

im sorry if i seemed like a bit dunce here, but i've been doing work for six straight hours and my brain is pretty drained right now. you actually did help me a lot

OpenStudy (driftracer305):

no.... no problem.... i understand........ it was good thou....... i hope u understood it....

OpenStudy (anonymous):

:)